OpenStudy (anonymous):

can anyone help me please!!?? For each polyhedron, use Euler’s Formula to find the missing number. 4. Faces: 9 Edges: Vertices: 14 6. Faces: Edges: 24 Vertices: 16

OpenStudy (amistre64):

i always have to try a square to remember it .... theres a 2 in there as well

OpenStudy (amistre64):

F+V=E+2 ... simple math

OpenStudy (anonymous):

so number 4 has 25 edges?

OpenStudy (amistre64):

9+14 = E + 2 9+12 + 2 = E + 2 E = 9+12, id say 21 edges